OpenStudy (jjuden):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = –2x^2 + x – 3

OpenStudy (jjuden):

a 1 point in common; vertex on x-axis b 2 points in common; vertex below x-axis c 2 points in common; vertex above x-axis d no points in common; vertex below x-axis

OpenStudy (sidsiddhartha):

first find the discriminant which is (b^2-4ac) here b=1,a=-2,c=-3

OpenStudy (jjuden):

its too late im done with this question

OpenStudy (sidsiddhartha):

if b^2 – 4ac > 0 then two real solutions which are different.................(1) if b^2 – 4ac = 0 then two real solutions which are equal.........................(2) if b^2 – 4ac < 0 then two complex solutions which are different..............(3) in ur problem (b^2-4ac)=1-24=-23 so no (3) is valid so ur ans is d no points in common; vertex below x-axis